Can a fluidized bed heat exchanger produce superheated steam using water and heat transfer fluid as inlets?
Answer #1

Hii
If we supply some amount of water and a fuel and the fuel is combusted then the temperature can easily reach above the boiling point of water.
We can support this with the help of a example -
One practical difficulty that you might be thinking is the material of construction of the packing and the packed bed column.
Packed bed is made up of material like asbestos and other refractory materials that can withatand a temperature as high as even more than 1000 o C.
Let us understand with a example
Let us say that the feed is 1 mole of water and one mole of
Propane,and we are supplying a excess of 100 % of oxygen.
Let us see the combustion reaction first

C3H8+5O2=3CO2+4H2O

Amount of oxygen required for the complete combustion = 5 moles
Since the amount of oxygen supplied is 100 percent excess amount of oxygen will be 10 moles.

We can supply pure oxygen or air, let us assume that we are supplying air then moles of N2 will be

10*0.79/0.21=37.61 moles

All this amount will remain unreacted

Moles of N2 remaining = 37.61 moles

Amount of CO2 released =3 moles
Amount of H2O = 4 moles
Amount of O2 remaining =10 moles
Total moles remaining =54 moles

Heat capacity of all these are less than 50J/mole.K, so let us take average heat capacity as 30 J/mole. K

Let us assume that the reactants are at 298 K

heat of combustion of propane = - 2210 KJ/mole
Heat released here will be used for increasing temperature of remaining flue gas.

Let the final temperature be T
The heat absorbed =Total moles *heat capacity *(temperature change)
=54*50*(T-298) =heat released =221000
Since (1kJ=1000 J)
T-298=818
T=1116 K = 833 o C (T(K) =273+T (o C))
Although we have taken approximation but we can see that temperature will reach as high as 833 o C.
Even if some heat is lost temperature can easily reach 100 o C, which is required for producing superheated steam.

Thank you